OpenStudy (anonymous):

39. If a spherical ball is enlarged so that its surface area is 9 times greater than its original surface area, then the original radius was multiplied by _________.

OpenStudy (anonymous):

What is the formula for a sphere's surface area?

OpenStudy (anonymous):

4 x pi x r^2

OpenStudy (anonymous):

You mean 4 x pi x r^2 = a (I forget the letter that is used to represent surface area) anyway, if the area is 9 times as large then you get 9a. Solve for a

OpenStudy (anonymous):

I'm confused...

OpenStudy (anonymous):

All formula's have two sides. Since we know that the the area is 9 times as large we can multiply the area by 9, with me so far?

OpenStudy (anonymous):

don't just give them the answer, that's worse then not helping at all.

OpenStudy (anonymous):

No, these are practice problems I found on my own to study. It is very helpful. Thanks!

OpenStudy (anonymous):

4πr^2 = S 4π(x*r)^2 = 9S Keep going with it :)

OpenStudy (anonymous):

If you are trying to understand the problem, your understanding will not be improved by getting the answer given to you, or even just an equation, true understanding is being able to do the whole thing on your own.

OpenStudy (anonymous):

Ivory is right ;) My fault for giving it away so quickly